# Computational Complexity of Sufficiency in Decision Problems

### v1

@@ Abstract
All results are machine-checked in Lean 4 ($\sim$5,000 lines, 200+ theorems).
@@ 1.3 Machine-Checked Proofs
All results are formalized in Lean 4 [7] ($\sim$5,000 lines, 200+ theorems). The formalization verifies the reduction mappings and combinatorial lemmas; complexity class membership follows by composition with TAUTOLOGY and $\exists\forall$-SAT.
@@
We contribute (i) a reusable Lean 4 framework for polynomial-time reductions with explicit polynomial bounds; (ii) the first machine-checked coNP-completeness proof for a decision-theoretic sufficiency problem; and (iii) a complete complexity landscape for coordinate sufficiency under explicit encoding assumptions. Prior work studies decision complexity in general or feature selection hardness, but does not formalize these reductions or establish this landscape in Lean.
@@ 9.1 Computational Decision Theory
Closest to our contribution is the literature on feature selection and model selection hardness, which proves NP-hardness of selecting informative feature subsets and inapproximability for minimum feature sets [3, amaldi1998complexity]. Those results analyze predictive relevance or compression objectives. We study decision relevance and show coNP-completeness for sufficiency checking, a different quantifier structure with distinct proof techniques and a full hardness/tractability landscape under explicit encoding assumptions, mechanized in Lean 4. The formalization aspect is also novel: prior work establishes hardness on paper, while we provide machine-checked reductions with explicit polynomial bounds.
@@ 10 Conclusion
Proofs are machine-checked in Lean 4 (see Appendix A for proof listings). The formalization verifies the reduction mappings and equivalence theorems; complexity classifications follow from standard results (TAUTOLOGY is coNP-complete, $\exists\forall$-SAT is $\Sigma_{2}^{\textsf{P}}$-complete) under the encoding model of Section 2.4. In particular, the strengthened gadget showing that non-tautologies yield instances with all coordinates relevant (hence $k^{*}=\Omega(n)$) is also mechanized (see Appendix A).
@@ Appendix A Lean 4 Proof Listings
The complete Lean 4 formalization is included in the appendix and companion artifact (Zenodo DOI is listed on the paper title page).
The formalization is not a transcription; it exposes a reusable reduction library with compositional polynomial bounds, enabling later mechanized hardness proofs to reuse these components directly.
A.1 On the Nature of Definitional Proofs
The Lean proofs follow directly from the formal definitions and standard complexity-theoretic constructions. The mechanization provides machine-checked verification and precision.
Definitional vs. derivational proofs. The core theorems establish definitional properties and reduction constructions. For example, the polynomial reduction composition theorem (Theorem A.1) proves that composing two polynomial-time reductions yields a polynomial-time reduction. The proof follows from the definition of polynomial time: composing two polynomials yields a polynomial.
Precedent in complexity theory. This pattern occurs throughout classical complexity theory:
  • Cook-Levin Theorem (1971): SAT is NP-complete. The proof constructs a reduction from an arbitrary NP problem to SAT. The construction itself is straightforward (encode Turing machine computation as boolean formula), but the insight is recognizing that SAT captures all of NP.
  • Ladner’s Theorem (1975): If P $\neq$ NP, then NP-intermediate problems exist. The proof is a diagonal construction—conceptually simple once the right framework is identified.
  • Toda’s Theorem (1991): The polynomial hierarchy is contained in P#P. The proof uses counting arguments that are elegant but not technically complex. The profundity is in the connection between counting and the hierarchy.
Why simplicity indicates strength. A definitional theorem derived from precise formalization is stronger than an informal argument. When we prove that sufficiency checking is coNP-complete (Theorem 3.6), we are not saying “we tried many algorithms and they all failed.” We are saying something general: any algorithm solving sufficiency checking solves TAUTOLOGY, and vice versa. The proof is a reduction construction that follows from the problem definitions.
Where the insight lies. The semantic contribution of our formalization is:
@@ A.1 On the Nature of Definitional Proofs
    Precision forcing. Formalizing “coordinate sufficiency” in Lean requires stating exactly what it means for a coordinate subset to contain all decision-relevant information. This precision eliminates ambiguity about edge cases (what if projections differ only on irrelevant coordinates?).
@@
    Reduction correctness. The TAUTOLOGY reduction (Section 3) is machine-checked to preserve the decision structure. Complexity reductions are subtle and easy to mis-specify. We therefore mechanize the definitions and key reductions in Lean to rule out common formalization mistakes.
@@
    Encoding-regime separation. Theorem 4.1 contrasts explicit-state polynomial-time (polynomial in $|S|$) with a succinct worst-case ETH lower bound witnessed by a hard family with $k^{*}=n$ requiring $2^{\Omega(n)}$ time. Intermediate regimes are not ruled out by the lower-bound statement.
What machine-checking guarantees. The Lean compiler verifies that each proof step is valid and each definition is consistent; the mechanization does not introduce additional axioms beyond the chosen foundations. The mechanization contains no sorry placeholders. Reviewers can independently verify the proofs by running lake build in the mechanization directory and inspecting the companion artifact.
Comparison to informal complexity arguments. Prior work on model selection complexity (Chickering et al. [4], Teyssier & Koller [18]) presents compelling informal arguments but lacks machine-checked proofs. Our contribution is not new wisdom—the insight that model selection is hard is old. Our contribution is formalization: making “coordinate sufficiency” precise enough to mechanize, constructing verified reductions, and proving the complexity results hold for the formalized definitions.
This follows the tradition of verified complexity theory: just as Nipkow & Klein [11] formalized automata theory and Cook [6] formalized NP-completeness in proof assistants, we formalize decision-theoretic complexity. The proofs are simple because the formalization makes the structure clear. Simple proofs from precise definitions are the goal, not a limitation.
A.2 Module Structure
The formalization consists of 33 files organized as follows:
  • Basic.lean – Core definitions (DecisionProblem, CoordinateSet, Projection)
  • AlgorithmComplexity.lean – Complexity definitions (polynomial time, reductions)
  • PolynomialReduction.lean – Polynomial reduction composition (Theorem A.1)
  • Reduction.lean – TAUTOLOGY reduction for sufficiency checking
  • Reduction_AllCoords.lean – Strengthened gadget: if $\varphi$ is not a tautology then every coordinate is relevant
  • Hardness/ – Counting complexity and approximation barriers
  • Tractability/ – Bounded actions, separable utilities, tree structure, FPT
  • Economics/ – Value of information and elicitation connections
  • Dichotomy.lean and ComplexityMain.lean – Summary results
  • HardnessDistribution.lean – Simplicity Tax Theorem (Section 8)
A.3 Key Theorems
Theorem A.1 (Polynomial Composition, Lean).
Polynomial-time reductions compose to polynomial-time reductions.
theorem PolyReduction.comp_exists
    (f : PolyReduction A B) (g : PolyReduction B C) :
    exists h : PolyReduction A C,
      forall a, h.reduce a = g.reduce (f.reduce a)
Theorem A.2 (Simplicity Tax Conservation, Lean).
The simplicity tax plus covered axes equals required axes (partition).
@@ A.3 Key Theorems
Theorem A.3 (Simplicity Preference Fallacy, Lean).
Incomplete tools always cost more than complete tools for $n>0$ use sites.
@@
A.4 Verification Status
The mechanization and its verification status are documented in the companion artifact; reviewers may consult the artifact for exact counts and files. The proofs compile and can be verified via the provided build instructions.

### v2

@@ Abstract
The reduction constructions and key equivalence theorems are machine-checked in Lean 4 ($\sim$5,000 lines, 200+ theorems); complexity classifications follow by composition with standard results (see Appendix A).
@@ 1.3 Machine-Checked Proofs
The reduction constructions and key equivalence theorems are machine-checked in Lean 4 [6] ($\sim$5,000 lines, 200+ theorems). The formalization verifies that the TAUTOLOGY reduction correctly maps tautologies to sufficient coordinate sets. Complexity class membership (coNP-completeness, $\Sigma_{2}^{P}$-completeness) follows by composition with standard complexity-theoretic results.
@@
We contribute (i) formal definitions of decision-theoretic sufficiency in Lean; (ii) machine-checked proofs of reduction correctness for the TAUTOLOGY and $\exists\forall$-SAT reductions; and (iii) a complete complexity landscape for coordinate sufficiency with explicit encoding assumptions. Prior work establishes hardness informally; we formalize the constructions.
@@ 9.1 Computational Decision Theory
Closest to our contribution is the literature on feature selection and model selection hardness, which proves NP-hardness of selecting informative feature subsets and inapproximability for minimum feature sets [4, 2]. Those results analyze predictive relevance or compression objectives. We study decision relevance and show coNP-completeness for sufficiency checking, a different quantifier structure with distinct proof techniques and a full hardness/tractability landscape under explicit encoding assumptions, mechanized in Lean 4. The formalization aspect is also novel: prior work establishes hardness on paper, while we provide machine-checked reductions with explicit polynomial bounds.
@@ 10 Conclusion
The reduction constructions and key equivalence theorems are machine-checked in Lean 4 (see Appendix A for proof listings). The formalization verifies that the TAUTOLOGY reduction correctly maps tautologies to sufficient coordinate sets; complexity classifications (coNP-completeness, $\Sigma_{2}^{\textsf{P}}$-completeness) follow by composition with standard complexity-theoretic results (TAUTOLOGY is coNP-complete, $\exists\forall$-SAT is $\Sigma_{2}^{\textsf{P}}$-complete). The strengthened gadget showing that non-tautologies yield instances with all coordinates relevant is also formalized.
@@ Appendix A Lean 4 Proof Listings
The complete Lean 4 formalization is available in the companion artifact (Zenodo DOI listed on the title page). The mechanization consists of approximately 5,000 lines across 33 files.
A.1 What Is Machine-Checked
The Lean formalization establishes:
@@ A.1 What Is Machine-Checked
    Correctness of the TAUTOLOGY reduction: The theorem tautology_iff_sufficient proves that the mapping from Boolean formulas to decision problems preserves the decision structure (accept iff the formula is a tautology).
@@
    Decision problem definitions: Formal definitions of sufficiency, optimality, and the decision quotient.
@@
    Economic theorems: The Simplicity Tax conservation laws and hardness distribution results.
Complexity classifications (coNP-completeness, $\Sigma_{2}^{P}$-completeness) follow from informal composition with standard results (TAUTOLOGY is coNP-complete, etc.). The Lean proofs verify the reduction constructions; the complexity class membership is derived by combining these with established theorems from complexity theory.
A.2 Polynomial-Time Reduction Definition
We use Mathlib’s Turing machine framework to define polynomial-time computability:
/-- Polynomial-time computable function using Turing machines -/
def PolyTime {  : Type} (ea : FinEncoding ) (eb : FinEncoding )
    (f :   ) : Prop :=
  Nonempty (Turing.TM2ComputableInPolyTime ea eb f)
/-- Polynomial-time many-one (Karp) reduction -/
def ManyOneReducesPoly {  : Type} (ea : FinEncoding ) (eb : FinEncoding )
    (A : Set ) (B : Set ) : Prop :=
   f :   , PolyTime ea eb f   x, x  A  f x  B
This uses the standard definition: a reduction is polynomial-time computable via Turing machines and preserves membership.
A.3 The Main Reduction Theorem
Theorem A.1 (TAUTOLOGY Reduction Correctness, Lean).
The reduction from TAUTOLOGY to SUFFICIENCY-CHECK is correct:
theorem tautology_iff_sufficient ( : Formula n) :
    .isTautology  (reductionProblem ).isSufficient Finset.empty
This theorem is proven by showing both directions:
  • If $\varphi$ is a tautology, then the empty coordinate set is sufficient
  • If the empty coordinate set is sufficient, then $\varphi$ is a tautology
The proof verifies that the utility construction in reductionProblem creates the appropriate decision structure where:
  • At reference states, accept is optimal with utility 1
  • At assignment states, accept is optimal iff $\varphi(a)=\text{true}$
A.4 Economic Results
The hardness distribution theorems (Section 8) are fully formalized:
@@ A.4 Economic Results
A.5 Module Structure
  • Basic.lean – Core definitions (DecisionProblem, sufficiency, optimality)
  • Sufficiency.lean – Sufficiency checking algorithms and properties
  • Reduction.lean – TAUTOLOGY reduction construction and correctness
  • Complexity.lean – Polynomial-time reduction definitions using mathlib
  • HardnessDistribution.lean – Simplicity Tax theorems
  • Tractability/ – Bounded actions, separable utilities, tree structure
A.6 Verification
The proofs compile with Lean 4 and contain no sorry placeholders. Run lake build in the proof directory to verify.